could you give some points about trig. I am always confused with trig about how to calculate cos46, tan32, sin54 withot a calculator..... could you pls clear this doubt

Grade:8

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
Try relating them with common values
sin0,30,45,60,90
for eg for sin15 use sin30 value
for sin54 we do like this
90-2*18=3*18
so
sin2*18=sin3*18
using sin2A and sin3A we can find sin18
then again we get sin3*18 by sin3A
similary cos36=sin54
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty